# DAWN: Designing Distributed Agents in a Worldwide Network

### v2

@@ III-A Principal Agent
In many agentic implementations, the action space or toolset is limited to the local resource repository. This is typically due to three factors: 1) specialized agents naturally operate within a limited action space[42, 9, 43], 2) the context window of LLMs restricts the number of resources they can handle, and 3) if the agent has more than 20 tools to choose from, the accuracy in selecting the right tool drops as suggested by OpenAI[44]. (Our own experimentation found this number to be less than 10.) However, fully autonomous agents need to navigate real-world problems where the action space is vast, and numerous resources are available. DAWN’s rationale to use distributed Gateway Agents follows directly from this tenet. For example, if the Principal Agent needs to perform an action beyond an LLM’s capabilities—such as online search, booking a ticket, or writing code—it will request the relevant resource (e.g., a search tool, booking agent, or coding agent) from Gateway Agents. This flexibility enhances the Principal Agent’s autonomy and allows it to handle a broader range of use cases. However, this flexibility of this type comes with a cost. A Principal agent might get overlapping, conflicting, malicious or faulty resources. Dealing with these issues in software engineering will be quite a challenge. Overlapping resources require special handling to determine which resource is best in terms of functional as well as non-functional performance characteristics. In the case of malicious agents are introduced, whether inadvertent or otherwise, these agentic resources will render the entire distributed agentic application vulnerable to attack or worse, loss of faith. This is precisely why the safety and security of such applications are of crucial importance. Such faulty resources are an inherent part of the internet as well and people have been working to keep the defense line strong. This is the price we must pay for an open, participatory system. The alternate is a closed one and the trust that one provider will make it safe, secure, but inflexible.
@@ III-B Gateway Agent
The Gateway Agent plays an essential role in connecting globally deployed resources (tools, agents, and agentic applications) with the Principal Agent. (There is a trivial case of human operators acting as the Principal Agent and using Gateway Agents for additional services under the No-LLM mode but the LLM-based applications are the central focus of this paper.) The Gateway Agent offers several services of which two are primary: a registry where resources are registered and an intelligent match-and-retrieve service for searching relevant resources given a query or task list. When the Principal Agent sends queries describing its need in natural language to the Gateway Agents, the Gateway Agents in turn search their registries and return the most appropriate resources to the Principal Agent. Each Gateway Agent may apply various filters and guardrails to select and trim the number of resources returned. Security measures such as access control and resource testing are integrated to reinforce safety of the system.
@@
When a resource is first registered with a Gateway Agent, the Gateway Agent has only limited information about it. It is screened for safety, security, compliance and other readily visible characteristics. Over time, the Gateway Agent also gathers utilitarian characteristics such as P50/P90 latency, cost, completion rate, etc. The registry may also include metrics such as success rate, average execution time, availability, and load capacity. These metrics allow the orchestrator to understand and optimize resource utilization and is then used to score, rate and select the best resource for a given user function.
Figure 4: Principal Agent and Gateway Agents workflow. For each user request, the Principal Agent first plans, produces a task list, and then requests resources from the Gateway Agents. The Gateway Agents return the most suitable resources to the Principal Agent, which is then used to first compose, and then orchestrate and execute the task list.
@@
Research has shown that Large Language Models (LLMs) can automatically communicate with RESTful APIs to execute tasks through them. RestGPT[45], for example, showcases this potential by leveraging the coding capabilities of LLMs to generate parsing code based on structured schemas. This approach automates both the generation of API calls and the parsing of responses and furthermore the LLM acts as a backup parser when necessary.
@@
If a resource provider fails to provide validation, the Gateway Agent will temporarily suspend the resource until the required tests are successfully completed, and the validation is provided. Likewise, resources managed by the Gateway Agents will undergo a policy check and alignment with guidelines through empirical testing and automated review [open ai store link]. If a resource fails to pass this validation process, indicating discrepancies between its description and actual behavior, the Gateway Agents will disqualify the agent from participating until a more accurate description is generated.
@@
Resource Retrieval: A Gateway Agent handles queries from the Principal Agent, and searches its registry for the most relevant resources. Similar to traditional information retrieval methods [50], resource descriptions and examples may be treated like documents. Semantic search [51] may be used to identify the most relevant resources by converting both the queries and resource descriptions into vector embeddings and then performing similarity searches based on those embeddings. To further improve search accuracy, additional techniques like attribute filtering and keyword search are incorporated. Moreover, fine-tuning LLMs on query-resource pairs is another viable approach to enhance search results [52, 41]. While DAWN imposes some base functional and non-functional requirements on Gateway Agents along with stringent safety and security guardrails, its internals are left to the discretion of the Gateway owner. Some Gateway Agents may only have traditional software and classical data repositories, while others may be fully functioning LLM-driven Agents.
@@
Multiple Gateway Agents: Organizations and developers around the globe will build and manage their own Gateway Agents following an open protocol. Each Gateway Agent offers unique resources proprietary to the entity in question (organization or developer). At any given time, the Principal Agent has a list of Gateway Agents it is connected with. The Principal Agent submits tasks and user intents to the listed Gateway Agents who then return the best resources available that may address all or a subset of the tasks or subtasks associated with the user request. The Principal Agent collects the search results and selects the highest-ranked (and most appropriate for the task) resources from them. This selection is based on factors like relevance, latency, cost, reputation, prior use, success and failure rates, and other metadata collected by the Gateway Agents about these agents over time. These metadata include information such as rating, latency, and performance of the registered agents. If the returned resources serve to complete the user tasks, then the Principal Agent’s task list is complete. If not, the cycle continues directed by the Principal Agent.
@@
Gateway Agents may disconnect for maintenance or upgrades when idle or if its pending requests can be queued. Gateway Agents designed with memory and context mechanism to hold unaddressed user requests in a queue are able to connect and disconnect on their own schedules. A new Gateway Agent may join the Principal Agent’s list at any time but new Gateway Agents may only participate in a subsequent user request from the Principal Agent. This stands to reason as it might not have the context to address an ongoing task list. The Principal Agent continuously evaluates Gateway Agents based on the quality of the resources they provide. This evaluation helps assign weights and ratings to Gateway Agents and the resources they offer. Furthermore, users have the option to rate and assess the final execution of their request, including how each element was handled. These user ratings may also impact the ratings of Gateway Agents. The Principal Agent also uses its own memory module to learn from its own experiences. Fig. 4 illustrates the workflow between a Principal Agent and a Gateway Agents in full autonomous mode, incorporating most of the functionalities discussed above.
@@ III-C Orchestration Layer
The Orchestration Layer in the Principal Agent, serves as a central component that enables the system’s versatility by supporting different operational modes, including No-LLM, Copilot, and hybrid modes that combine No-LLM, Copilot, and LLM Agent functionalities. In this hybrid mode, a human operator defines a workflow as an execution graph, where each node corresponds to a subtask, and the most suitable resource for each subtask is selected from those returned by the Gateway Agents. The Orchestration Layer is responsible for enforcing the rules, logic, and sequencing necessary to ensure that each subtask is executed in the correct order, adhering to the predefined workflow structure.
Additionally, the Orchestration Layer maintains the global state, tracking the progress of task execution across all nodes in the workflow and across all the participating Gateway Agents and global agentic resources. This global state ensures that each resource has the contextual information needed to perform its assigned subtask effectively. By maintaining and updating this global state, the system ensures that subtasks are executed with full awareness of prior actions and the broader context of the overall task.
@@ III-D Communication Layer
The Communication Layer is responsible for managing the flow of information between various system components and resources, ensuring seamless interaction across the platform. It handles the messaging protocols that ensure messages are properly parsed, interpreted, and routed to their correct destinations. Two primary types of messages are exchanged among the system components and resources.
The first type is the query sent by the Principal Agent to the Gateway Agent. The payload initially contains the subtasks and their descriptions. This step also involves data sanitization and purposeful query generation to craft queries and tasks in a way that focuses on the specific information or task needed, without providing excessive user details. During task execution, if a resource fails, the Principal Agent will reach out to the Gateway Agents. The Principal Agent will provide the current subtask and its description as well as the completed subtasks and interaction history. This additional context enables the Gateway Agent to make informed decisions about which new resources to retrieve. In response, the Gateway Agent then returns a list of suitable resources along with manifest files that contain the full record of the functional and non-functional capabilities of the resources.
The second type of message is the execution command sent by the Principal Agent to the Gateway Agents’ resource executors. These message contain the necessary execution details, such as the resource endpoint and input parameters required for the task. A given resource executor may then respond with the result of the execution. In the event of an error during execution, the message will include an error notification to inform the Principal Agent, allowing for appropriate error handling and adjustments.
@@ III-E Context Layer
The Context Layer is vital for enabling agentic operations by managing memory and task-related context. It provides three key components for memory management: a scratchpad to track the agent’s internal reasoning process, a message pool to record communication between the Principal Agent and the Gateway Agent, and a memory bank to store long-term data and user preferences, supporting long context, in-session, cross-session, multi-turn conversations resulting in personalized user experience. The Context Layer dynamically assembles prompts using data from both the message pool and memory bank, ensuring that the Principal Agent is aware of the overall task, task progress, and available resources.
Figure 5: Possible attacks on agentic networks and how DAWN mitigates such attacks. The Principal Agent and each of the Gateway Agents implement safety, security and compliance screening and filter to scrub agentic resources inputs and outputs.
@@ III-F Security, Safety, and Compliance Layer
The Security, Safety, and Compliance Layer establishes a robust security foundation for the entire framework, addressing the safety, security, and compliance needs of the system. All frameworks and platforms that employ distributed resources must implement robust security systems to avoid risks that may be introduced when unsecured resources are connected to an enterprise’s gateway agent. While extensive research and practices have been published on traditional common security issues such as authentication, access control, privacy, and policy enforcement [53, 54, 55], an agentic network equipped with LLMs introduces additional security and privacy threats especially as resources worldwide begin collaborating. For LLM agents, two primary sources of threats exist.
@@
There is extensive research dedicated to addressing both the inherent vulnerabilities of LLMs and attacks on agentic networks[62] . To protect DAWN framework against the knowledge poisoning, selecting reputable and transparent resources is essential. These resources must provide clear data usage policies and strong security records [63]. Tools and agents should also undergo third-party testing and sanitization, as demonstrated by ToolEmu [64], which proposes a framework for emulating tool execution through a language model. This approach allows scalable testing of the resources across various toolsets. Defense against functional manipulation begins with proactive security measures. For example, when employing third-party LLM agents, users should limit data sharing, particularly of sensitive personal information (PII). DAWN supports dynamic configurations where a co-pilot model integrates LLM-based and deterministic workflows, ensuring privacy and enterprise compliance often enforced by guardrails through algorithms designed to filter inputs and outputs[65, 66]. Incorporating deterministic steps into the workflow allows the enforcement of organizational policies and ensures user compliance through a LookUp tool, addressing potential instances where LLMs may bypass established guidelines. Additionally, increasing the reasoning capabilities of LLM agents is essential to counteract output manipulation. In the DAWN framework, the impact of output manipulation attacks is reduced by augmenting the Principal Agent with reasoning capabilities and access to a memory layer. Additionally, the modular architecture design enables the tracing of agentic attacks. Fig. 5 shows the attacks specific to agentic networks and DAWN’s approach in mitigating them.
The DAWN framework leverages Cisco’s expertise in safety and security to eliminate network attacks through traditional measures. Additionally, the outlined steps provide multiple layers of security within our agentic network, ensuring the system’s integrity, trustworthiness, and regulatory compliance.
Figure 6: Application of DAWN in an Human Resources recruiting use case in Agent + Copilot + No-LLM mode
@@ IV Application of DAWN
The modular design of DAWN framework provides users with the flexibility to select the most appropriate working mode for their application, tailored to the specific needs of various use cases. As outlined in the introduction of this paper, the primary modes include No-LLM, Copilot, and LLM Agent mode. For scenarios requiring high determinism, users manually define plans, select suitable resources through the Gateway Agent, and assemble a No-LLM application within the orchestration layer, ensuring precise and predictable outcomes. In situations where uncertainty is present and decisions must be made dynamically based on real-time feedback, users manually construct a decision graph, incorporating an LLM to intelligently select which subgraph to execute based on evolving conditions. For cases with flexible workflow, users delegate the entire task to the Principal Agent, which autonomously plans, finds resources from the Gateway Agent, and constructs agentic applications. This adaptability ensures that the framework can efficiently handle a wide range of complex and diverse tasks.
@@
To demonstrate DAWN’s capability, a walk-through example of a Human Resources (HR) recruiting use case is presented below. Finding the right candidate for a specific role in an organization involves multiple steps, most of which are repetitive and human-intensive, making them well-suited for LLM agents. When providing an HR application through DAWN, it is important to consider that the hiring process of each organization is optimized by HR experts according to that organization’s needs and policies. And it is crucial that not a single step be missed. Certain steps, such as offer negotiation with a candidate, require high security and are best handled by humans. Therefore, customers using DAWN for their agentic application are advised to use the framework in a hybrid mode.
Fig. 6 illustrates an HR example in a hybrid mode of Agent + Copilot + No-LLM. This hybrid mode includes a copilot that allows the customer to design the workflow manually, ensuring that none of the essential steps in the hiring process are neglected. The workflow is a directed graph designed by the customer, with most steps of the hiring process assigned to agentic nodes. These nodes leverage the full autonomous agent mode, where the Principal Agent requests resources from a Gateway Agent. In full autonomous mode, multiple specialized agents are seamlessly coordinated by the Principal Agent’s orchestration and the Gateway Agent’s resource retrieval. This setup promotes smooth cooperation and task efficiency. Additionally, the adaptive workflow allows the Principal Agent to dynamically adjust team members according to task progress. As it advances, the Principal Agent continuously communicates with the Gateway Agent for further assistance, ensuring the plan is regularly evaluated and updated if needed. This approach also supports iterative and traceable problem-solving, ensuring transparency and accountability throughout the entire process, making complex task management more effective.

### v3

@@ III-A Principal Agent
For example, if a Principal Agent needs to perform an action beyond an LLM’s capabilities – such as online search, booking a ticket, or writing code – it will request the relevant resource (e.g., a search tool, booking agent, or coding agent) from Gateway Agents. This flexibility enhances the Principal Agent’s autonomy and allows it to handle a broader range of use cases. However, flexibility of this type comes with a cost. The Principal agent might get overlapping, conflicting, malicious or faulty resources. Overlapping or duplicate resources require special handling to determine which resource is best in terms of functional as well as non-functional performance characteristics. In case malicious agents are introduced, whether inadvertently or otherwise, these agentic resources will render the entire distributed agentic application vulnerable to attack and worse, loss of faith. This is precisely why the safety and security of such applications are of crucial importance. Such faulty resources are an inherent part of the internet as well and people have been working to keep the defense line strong. This is the price we must pay for an open, participatory system. The alternate is a closed one and the trust that one provider will make it safe, secure, but inflexible.
@@ III-B Gateway Agent
The Gateway Agent plays an essential role in connecting globally deployed resources (tools, agents, and agentic applications) with the Principal Agent. The Gateway Agent offers several services of which two are primary: a registry where resources are registered and an intelligent match-and-retrieve service for searching relevant resources given a query or task list. When the Principal Agent sends queries describing its need in natural language to the Gateway Agents, the Gateway Agents in turn search their registries and return the most appropriate resources to the Principal Agent. Each Gateway Agent may apply various filters and guardrails to select and trim the number of resources returned. Security measures such as access control and resource testing are integrated to reinforce safety of the system.
@@
When a resource is first registered with a Gateway Agent, the Gateway Agent has only limited information about it. It is screened for safety, security, compliance and other readily visible and externalized characteristics. Over time, the Gateway Agent gathers utilitarian characteristics such as P50/P90 latency, cost, completion rate, prior use, etc. The registry may also include metrics such as success rate, average execution time, availability, and load capacity. These metrics allow the orchestrator to optimize resource utilization and then use it to score, rate and select the best resource for a given user function.
Figure 4: Principal Agent and Gateway Agents workflow. When invoked with an intent for an application or solution, the Principal Agent first plans, produces a task list, and then requests resources from the Gateway Agents. The Gateway Agents return the most suitable resources to the Principal Agent, which is then used to first compose, then orchestrate, and next, execute the task list.
@@
LLMs can automatically communicate with RESTful APIs to execute tasks through them. RestGPT[46], for example, showcases this potential by leveraging the coding capabilities of LLMs to generate parsing code based on structured schemas. This approach automates both the generation of API calls and the parsing of responses and furthermore the LLM acts as a backup parser when necessary.
@@
If a resource provider fails to provide validation, the Gateway Agent will temporarily suspend the resource until the required tests are successfully completed, and the validation is provided. Likewise, resources managed by the Gateway Agents will undergo a policy check and alignment with guidelines through empirical testing and automated review. If a resource fails to pass this validation process, indicating discrepancies between its description and actual behavior, the Gateway Agents will disqualify the agent from participating until a more accurate description is generated.
@@
Resource Retrieval: A Gateway Agent handles queries from the Principal Agent, and searches its registry for the most relevant resources. Similar to traditional information retrieval methods [51], resource descriptions and examples may be treated like documents. Semantic search [52] may be used to identify the most relevant resources by converting both the queries and resource descriptions into vector embeddings and then performing similarity searches based on those embeddings. To further improve search accuracy, additional techniques like attribute filtering and keyword search are incorporated. Moreover, fine-tuning LLMs on query-resource pairs is another viable approach to enhance search results [53, 41]. While DAWN imposes some base functional and non-functional requirements on Gateway Agents along with stringent safety and security guardrails, its internals are left to the discretion of the Gateway Agent’s owner. Some Gateway Agents, however, may only have traditional software and classical data repositories.
@@
Multiple Gateway Agents: Organizations and developers around the globe will build and manage their own Gateway Agents. While DAWN prescribes the components and services included in a Gateway Agent, the implementation is left to the organizations that build and manage them. To participate in a agentic solution with a Principal Agent, the Gateway Agent must have an open protocol. Each Gateway Agent offers unique resources proprietary to the entity in question (organization and its developers). Gateway Agents are also by design hybrid entities combining classical and agentic components to cover both AI and non-AI functionality. At any given time, the Principal Agent has a list of Gateway Agents it is connected with. These are loosely assembled entities along the lines of cloud-backed APIs. The Principal Agent submits tasks and user intents to the connected Gateway Agents which then return the best resources available that may address all or a subset of the tasks or subtasks associated with the user request. The Principal Agent collects the results and selects the highest-ranked (and most appropriate for the task) resources from them. This selection is based on factors like relevance, latency, cost, reputation, prior use, success and failure rates, and other metadata collected by the Gateway Agents about these agents over time. Gateway Agents retain and learn similar data about agents attached to themselves and also which services are requested and used by Principal Agents. This metadata may include information such as rating, latency, and performance of the registered agents. If the returned resources serve to complete the user tasks, then the Principal Agent’s task list is complete. If not, the cycle continues as directed by the Principal Agent.
@@
Gateway Agents may disconnect for maintenance or upgrades when idle or if its pending requests can be queued. Gateway Agents designed with memory and context mechanism to hold unaddressed user requests in a queue are able to connect and disconnect on their own schedules. A new Gateway Agent may join the Principal Agent’s list at any time but new Gateway Agents may only participate in a subsequent user request from the Principal Agent. This stands to reason as it might not have the context to address an ongoing task list. The Principal Agent continuously evaluates Gateway Agents based on the quality of the resources they provide. This evaluation helps assign weights and ratings to Gateway Agents and the resources they offer. Furthermore, end-users and developers have the option to rate and assess the final execution of their request, including how each element was handled. These user ratings may also impact the ratings of Gateway Agents. The Principal Agent also uses its own memory module to learn from its own experiences. Fig. 4 illustrates the workflow involving a Principal Agent and two Gateway Agents, incorporating most of the functionalities discussed above.
@@ III-C Orchestration Layer
The Orchestration Layer enables the system’s versatility by supporting different operational modes, including No-LLM, Copilot, and LLM agent modes. These services enable enable multi-tenant and multi-user concurrency and keep each operating application distinct and separate from all the others. When one such application is composed by the human user or the Principal agent, the Orchestration Layer services are responsible for enforcing the rules, logic, and sequencing necessary to ensure that each subtask is executed in the correct order. Internally, the task graph created from the each task list is implemented as a finite state machine whose nodes are agents or agent ensembles. Additionally, the Orchestration Layer maintains the global state, tracking the progress of task execution across all nodes in the workflow and across all the participating Gateway Agents (and global agentic resources). This global state ensures that each resource has the contextual information needed to perform its assigned subtask effectively. By maintaining and updating this global state, the system ensures that subtasks are executed with full awareness of prior actions and the broader context of the overall task.
@@ III-D Communication Layer
The Communication Layer is responsible for managing the flow of information between various system components and resources, ensuring seamless interaction across the platform. Its inherent services handle the protocols that ensure messages are properly parsed, interpreted, and routed to their correct destinations. Two primary types of messages are exchanged among the system components and resources.
The first type are the messages exchanged between the Principal Agent and its connected Gateway Agents. The payload from the Principal Agent contains the subtasks and their descriptions. This may involve data cleansing and purposeful query generation that focuses on the specific information or task needed, without providing excessive user details. During task execution, if a resource fails, the Principal Agent will reach out to the Gateway Agents to either repeat instructions or gather logs and metrics. The Principal Agent will provide the current subtask and its description as well as the completed subtasks and interaction history. This additional context enables Gateway Agents to make informed decisions about which new resources to retrieve. In response, the Gateway Agent then returns a list of suitable resources along with manifest files that contain the full record of the functional and non-functional capabilities of the resources. All of these interactions happen over channels set up by the Communication Layer.
The second type of message is the execution command sent by the Principal Agent to the Gateway Agents’ resource executors. These messages contain the necessary execution details, such as the resource endpoint and input parameters required for the task. A given resource executor may then respond with the result of the execution. In the event of an error during execution, the message will include an error notification to inform the Principal Agent, allowing for appropriate error handling and adjustments.
@@ III-E Context Layer
The Context Layer is vital for enabling agentic operations by managing memory and task-related context. It provides three key components for memory management: a scratchpad to track the agent’s internal reasoning process, a message pool to record communications between the Principal Agent and Gateway Agents, and a memory bank to store long-term data and user preferences, supporting long context, in-session, cross-session, multi-turn conversations resulting in personalized user experience. The Context Layer services dynamically assemble prompts using data from both the message pool and memory bank, ensuring that the Principal Agent is aware of the overall task, task progress, and available resources for all tasks and users that may be using the system at any given time.
Figure 5: Possible attacks on agentic networks and how DAWN mitigates such attacks. The Principal Agent and each of the Gateway Agents implement safety, security, and compliance screening to scrub agentic resources’ inputs and outputs.
@@ III-F Security, Safety, and Compliance Layer
The Security, Safety, and Compliance Layer establishes a robust security foundation for the entire framework, addressing the safety, security, and compliance needs of the system. All frameworks and platforms that employ distributed resources must implement robust security systems to avoid risks that may be introduced when unsecured resources are connected to an enterprise’s gateway agent. While extensive research and practices have been published on traditional common security issues such as authentication, access control, privacy, and policy enforcement [54, 55, 56], an agentic network equipped with LLMs introduces additional security and privacy threats especially as resources worldwide begin collaborating [23, 57]. The DAWN framework leverages available safety and security components and services to eliminate network attacks through traditional measures. Additionally, the outlined steps provide multiple layers of security within the agentic network, ensuring the system’s integrity, trustworthiness, and regulatory compliance. For LLM agents, two primary sources of threats exist.
@@
There is extensive research dedicated to addressing both the inherent vulnerabilities of LLMs and attacks on agentic networks[64] . To protect the DAWN framework against the knowledge poisoning, selecting reputable and transparent resources is essential. These resources must provide clear data usage policies and strong security records [65]. Tools and agents should also undergo third-party testing and sanitization, as demonstrated by ToolEmu [66], which proposes a framework for emulating tool execution through a language model. This approach allows scalable testing of the resources across various tool-sets. Defense against functional manipulation begins with proactive security measures. For example, when employing third-party LLM agents, users should limit data sharing, particularly of sensitive personal information (PII). DAWN supports dynamic configurations, particularly the co-pilot mode, that integrates LLM-based and deterministic workflows. This ensures privacy and enterprise compliance that are enforced by guardrails through algorithms designed to filter inputs and outputs[67, 68]. Incorporating deterministic steps into the workflow allows the enforcement of organizational policies and ensures user compliance through a LookUp tool, addressing potential instances where LLMs may bypass established guidelines. Additionally, increasing the reasoning capabilities of LLM agents is essential to counteract output manipulation. In the DAWN framework, the impact of output manipulation attacks is reduced by augmenting the Principal Agent with reasoning capabilities and access to a memory layer. Additionally, the modular architecture design enables the tracing of agentic attacks. Fig. 5 shows the attacks specific to agentic networks and DAWN’s approach in mitigating them.
Figure 6: Application of DAWN in an Human Resources recruiting use case in Agent, Copilot, and No-LLM modes
@@ IV Application of DAWN
The modular design of DAWN framework provides users with the flexibility to select the most appropriate working mode for their application, tailored to the specific needs of various use cases. As outlined in the introduction of this paper, the primary modes include No-LLM, Copilot, and LLM Agent mode. For scenarios requiring high determinism, users manually define plans, select suitable resources through the Gateway Agent, and assemble a No-LLM application within the orchestration layer, ensuring precise and predictable outcomes. In situations where uncertainty is present and decisions must be made dynamically based on real-time feedback, users manually construct a decision graph, incorporating an LLM to intelligently select which subgraph to execute based on evolving conditions. For cases with flexible workflows, users delegate the entire task to the Principal Agent, which autonomously, or along with human oversight, plans, finds resources from Gateway Agents, and constructs agentic applications. This adaptability ensures that the framework can efficiently handle a wide range of complex and diverse tasks.
@@
To demonstrate DAWN’s capability, a walk-through of the implementation of a Human Resources (HR) recruiting use case is presented below. Finding the right candidate for a specific role in any organization involves multiple steps that span weeks. Many of the steps are repetitive and therefore well-suited for agentic automation. They also require supervision and oversight at every stage. Most organizations have hiring processes in place and agentic software, such as this HR implementation of DAWN, needs to be built jointly with HR experts and according to that organization’s policies and guidelines. Certain steps, such as offer negotiation with a candidate, require high security and sensitivity and are best handled by humans. Therefore, developers and users are advised to apply DAWN in copilot mode.
Fig. 6 illustrates an HR example, a composite of Agent, Copilot, and No-LLM modes. This approach allows the Principal Agent to devise a plan once it is provided the requirements as instructions. The human then ensures that none of the essential steps in the hiring process is neglected. The supervision also ascertains that all of outputs that demand inspection are closely monitored. The workflow is a directed graph with most of the steps assigned to agentic nodes. These nodes leverage the autonomous agent mode where possible with the Principal Agent requesting resources from Gateway Agents. In fully autonomous mode, multiple specialized agents will be seamlessly coordinated by the Principal Agent’s orchestration and the Gateway Agent’s resource retrieval. However, here the deployment was supervised by an human. The agents and process flow deployed in this use cases were designed using the DAWN architecture and implemented using agents created using the LangGraph framework. Planning and composing was done using a Principal Agent and the agents themselves were registered with two different internal Gateway Agents. (While two were not strictly necessary, they illustrate how specialized agents may be registered with Gateway Agents with different owners.) The details of these agents are as follows.
  1. 1.
    The hiring process begins with creating a job description or JD. JDs have specific characteristics, mandatory sections, and descriptions. For this implementation, the team developed an agent to write the JD for a machine learning engineer using descriptions gathered from the web. The agent was provided with sample JDs. This step was developed in agentic mode with developers (proxies for HR experts in this case) reviewing the output and iteratively modifying the prompts until satisfied. The DAWN framework was used for iterative development and testing, and its intrinsic SSC layer was used to make sure the output JD was compliant to company, federal, and state policies and regulations. JDs are short documents and take only a few seconds of LLM time to process and produce new versions.
  2. 2.
    With a JD in place, the next agent in the workflow finds matching candidate profiles in LinkedIn using the LinkedIn API. The agent first pulls apart the different sections of the JD such as title, experience, education, skills, location, etc. It then applies these requirements to the LinkedIn advanced search API. Next, the LinkedIn filtering APIs are used to refine the results. Then, as a final step, the agent applies any organizational screens to further distill the results. This search-and-filter step is iterative and requires humans to review the number and quality of the retrieved profiles to adjust the agent’s input constraints. The agent functions autonomously but its output needed to be adjusted based on the job description and level of the intended hire. Searching and filtering takes but a few seconds but altogether this step can take hours because of the inspection required.
  3. 3.
    The third agent constructs the interview panel for each candidate for each job role. It uses the same pool of available interviewers and accesses their calendars to find empty slots. Scheduling itself is simple but more realistically contacting the candidate for available time slots is done by recruiters. In this implementation, only the calendars of the interviewers were checked. This agent interfaces with Microsoft Outlook to book these slots. No allowance was made for change requests though realistically that should also be accounted for. This agentic step was fully automated and took milliseconds to execute.
  4. 4.
    While agents can listen in to interviews, transcribe, make summaries, and conclusions, that is often impractical, socially unacceptable, and in some jurisdictions, outright illegal. In this implementation, the agent sends reminders to interviewers to document and upload the interview feedback within a time window, gathers and aggregates all the feedback from the database, and compiles that into a single document in preparation for the next phase - the interview debrief and decision meeting.
  5. 5.
    Once a decision is made, all the materials required for the hiring manager to make the final decision are done using existing tools and processes. This is the No-LLM ”agent” depicted in the figure. A lot of hiring experience and support from recruitment experts and financial advisors go into this step and it makes little sense to try and use AI except to gather information and make sure decisions adhere to policies and guidelines.
  6. 6.
    The last agent sets up onboarding education for the new employee. It compiles the briefing package that includes reading material, self-training, people to meet, orientation meetings to attend, and other information required of all newbies. And then are special sessions for those who may have specialized responsibilities. In this implementation this was conceived as a list to complete with appropriate calendaring. The agent sends reminders and calendar events after building a customized onboarding and orientation program for the new employee based on level and specialized function.
To measure the effectiveness of this DAWN architecture requires looking at success and failure rates in each component and each step of the workflow. Agentic software development covers the following agentic actions: plan, compose, discover, connect, orchestrate, observe, safeguard, and communicate. Therefore, each of these may be measured and evaluated, and taken together indicate how effective DAWN is. Of these, plan, and compose lie entirely within the Principal Agent’s purview. The remaining functions are distributed between the Principal and Gateway Agents. This paper has already addressed connect, orchestrate, observe, safeguard, and communicate to some extent. For the most part these functions mirror those used in traditional distributed systems and may be evaluated along the same lines. Plan and compose are novel in agentic software development. Therefore, in this experiment only the plan and compose elements of the process was evaluated. Planning involves producing a task list from the application intent while composing involves producing a task graph (finite state machine) from the task list and finding and selecting the most appropriate agents to perform the actions. The following algorithm was used to measure the success rate of the planner:
  1. 1.
    Provide the LLM with a fully qualified application intent and it outputs a task list. An LLM is prompted to generate 10 use case descriptions that demand the use of agentic AI, each with a task list and a graph of ”tasks that are not just sequential”. It produced 10 task lists and 10 task graphs. This serves as the ground truth.
  2. 2.
    Select a use case at random from the list and send it to the same LLM (as in step (1)) using the same prompt. Compare the task list against the ground truth.
  3. 3.
    Repeat step (2) 5 times. Measure repeatability in terms of all the tasks and the correct order of the tasks. (In some cases. reordering tasks may not matter and in these cases reordered tasks are marked correct provided they are complete in other respects.)
  4. 4.
    Optionally, repeat steps (2) and (3) with other LLMs
  5. 5.
    Compute the success rate.
The 10 LLM generated use cases each had 6-10 tasks (task descriptions each 1-2 sentences), some that required a sequential application of agents, and others, more complex graphs. The computation for steps (2) and (3) above was done manually and using a different LLM as a judge. The success rate for task list creation ranged from 70-85% – the lower number when a different LLM (OpenAI 4o-mini) was used for assessment from the original one used for task list creation. The complexity of the tasks varied quite a bit across the 10 use cases and that might also explain some of the variance in the results.
Composition requires finding the best matching agents for a given task. This is modeled as an information retrieval (IR) task. The IR task was setup as a separate experiment to measure task-to-agent matching accuracy against a well-known public IR dataset with 200 agents and over 20,000 queries. Given the agent registry with agent names, descriptions, and other metadata, a two-stage information retrieval system was developed to find the best fitting agent for each task. Stage 1 is a retriever backed by a vector database and stage 2 is an LLM-based re-ranker. The retriever produces a fixed list of top-5 agents for each task in the dataset and the re-ranker uses those candidates and any additional contextual information provided to generate the best candidate/s. Two metrics, Normalized Discounted Cumulative Gain (NDCG, measured at top k) and Recall (Recall@k), were computed. After stage 2, NDCG at k=1 and k=3 was 0.71 and 0.78, respectively, while Recall at k=1 and k=3 was 0.71 and 0.82, respectively.